OpenStudy (anonymous):

Solve the system by using a table. -0.2x-0.3y=0.2 0.4x-0.3y=1.4 The possible answers are; 1. (2,2) 2. (2,-2) 3. (2,1) 4. (-2,2) Please help I'm so confused:(

hero (hero):

You can solve it using a table, but you should first do something about those pesky decimals. Those are what may be causing your confusion. Simply multiply both sides of each equation by 10, then you'll have -2x - 3y = 2 4x - 3y = 14 Also you can express the system as 2x + 3y = -2 4x - 3y = 14 Now it looks something similar to what you are used to working with.

hero (hero):

Probably the quickest way to figure out which point is correct is to test each point by plugging it in to both equations. The point you choose has to work for both equations in order for it to be the correct one.

OpenStudy (anonymous):

Thank you @Hero !

hero (hero):

You're welcome

hero (hero):

@Briana_Brooke which answer choice did you pick?

OpenStudy (anonymous):

(-2,2) @Hero

hero (hero):

What about (2, -2)?

OpenStudy (anonymous):

oh! @Hero I must have gotten them switched around somehow

hero (hero):

That's okay

OpenStudy (anonymous):

can you help me with another one @Hero ?

hero (hero):

Close this question and post a new one.